Ionization potential of hydrogen Atom is 13.6eV.Hydrogen atoms in ground state are excited by monochromatic radiation of photon energy 12.1eV.According to Bohr`s theory, the spectral lines emitted by hydrogen will be.?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
Given Ionization Potential Energy = 13.6eV.
and also Photon Energy = 12.1eV
En=−13.6+12.1=1.5eVEn=−13.6+12.1=1.5eV
∴n^2=E1/En∴n2=E1/En
⇒13.6/1.5⇒13.6/1.5=9=9
⇒n=3⇒n=3
Thus transition occurs from 3rd3rd to 1st1st orbit.
∴∴ spectral lines emitted =∑Δn=∑Δn
⇒∑(3−1)⇒∑(3−1)
⇒∑(2)⇒∑(2)
⇒1+2=3
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)